# A simple controller design to achieve iso-damping robustness: Non-iterative data-driven approach based on fractional-order reference model

### v1

@@ 4.2 Example 2: Superiority over the conventional approach based on the reduced model
where $K_{fp},K_{fi},\lambda\in\mathbb{R}_{+}\cup\left\{0\right\}$. The tuning condition is listed in Table 4. Fig. 8 describes the initial input/output data obtained using $C_{FOPI}\left(z;\theta_{0}\right)$, where the simulation time for initial data collection is $2\,\mathrm{s}$ and the reference input $r_{[0:N]}^{D}$ is the unit step. Note that the initial data is collected on the basis of the actual plant $P\left(z\right)$ in (21).
@@
Table 4: Conditions for Example 2.
@@
Table 5: Tuning result of the proposed approach (Example 2).
@@
Table 6: Frequency domain property with the actual plant $P\left(z\right)$ (Example 2).
@@
Figure 8: Initial input/output data for the proposed controller tuning (Example 2).
@@
Figure 10: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{iso-m}\left(z;\theta^{\ast}\right)$, $P\left(z\right)C_{proposed}\left(z;\theta^{\ast}\right)$ (Example 2): (a) overview, and (b) enlarged phase plots around 12 Hz.
@@ 5 Discussion
In this study, we assume that the reference model is given by the designer. That is, the designer must give the appropriate reference model. An important future direction may be the simplification of the design procedure for the reference model. For example, the flat-phase property is frequency-domain specification, whereas it is more straightforward and intuitive to design the reference model in the time-domain. The Auto-DDC approach [7][27] may be effective for addressing this issue. One study has proposed the data-driven tuning of parameters of both the controller and the reference model [54]. In the future, we will examine the Auto-DDC strategy and the simultaneous tuning technique of the controller and the reference model in iso-IDFRIT.
6 Conclusion

### v2

@@ IV-B Example 2: Soft robot (superiority over the conventional model-based approach)
where $K_{fp},K_{fi},\lambda\in\mathbb{R}_{+}\cup\left\{0\right\}$. The tuning condition is listed in Table IV. Fig. 8 describes the initial input/output data obtained using $C_{FOPI}\left(z;\theta_{0}\right)$, where the simulation time for initial data collection is $2\,\mathrm{s}$ and the reference input $r_{[0:N]}^{D}$ is the unit step. Note that the initial data is collected on the basis of the actual plant $P\left(z\right)$ in (21).
@@
TABLE IV: Conditions for Example 2.
@@
TABLE V: Tuning result of the proposed approach (Example 2).
@@
TABLE VI: Frequency domain property with the actual plant $P\left(z\right)$ (Example 2).
@@
Figure 8: Initial input/output data for the proposed approach (Example 2).
@@
Figure 10: Bode plots of $L_{flat}\left(z\right)$, $P\left(z\right)C_{iso-m}\left(z;\theta^{\ast}\right)$, $P\left(z\right)C_{proposed}\left(z;\theta^{\ast}\right)$ (Example 2): (a) overview, and (b) enlarged phase plots around $12\,\mathrm{rad/s}$.
@@ V Discussion
In this study, we assume that the reference model is given by the designer. That is, the designer must give the appropriate reference model. An important future direction may be the simplification of the design procedure for the reference model. For example, the flat-phase property is frequency-domain specification, whereas it is more straightforward and intuitive to design the reference model in the time-domain. The Auto-DDC approach [62] may be effective for addressing this issue. One study has proposed the data-driven tuning of parameters of both the controller and the reference model [63]. In the future, we will examine the Auto-DDC strategy and the simultaneous tuning technique of the controller and the reference model in iso-IDFRIT.
VI Conclusion